OpenStudy (anonymous):

The amount $5.40 is 9% of what price?

OpenStudy (danjs):

\[5.40 = \frac{ 9}{ 100 }*x\]

OpenStudy (danjs):

5.40 is [=] 9% [9/100] of [*] what price [x]

OpenStudy (danjs):

\[100 * 5.40 = 9x\]

OpenStudy (anonymous):

?

OpenStudy (danjs):

x = 60

OpenStudy (agl202):

yes thats correct danjs

OpenStudy (danjs):

Just interpret the words into an equation is - = of - multiply what percent = x/100 a number - x